Emma and Mary are the two tallest basketball players on their high school team. Emma is 67 1/2 inches tall and Mary is 70 1/4 inches tall. How many inches taller is Mary than Emma
Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

Subtract Emma's height from Mary's height, then you'll know.
---------
It's arithmetic.

Answer by rfer(16322)

70 1/4 - 67 1/2=
70 1/4 - 67 2/4=
69 5/4 - 67 2/4= 2 3/4 inches